In a random sample of 14 residents of the state of Montana, the mean waste recycled per person per day was 1.4 pounds with a standard deviation of 0.71 pounds. Determine the 99% confidence interval for the mean waste recycled per person per day for the population of Montana. Assume the population is approximately normal. Step 1 of 2 : Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

Given that,

\bar x = 1.4

s = 0.71

n = 14

Degrees of freedom = df = n - 1 = 14 - 1 = 13

At 99% confidence level the t is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

critical value t\alpha /2  df = t0.005,13 =3.012    ( using student t table)

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 3.012 * (0.71 / \sqrt 14) = 0.57

The 99% confidence interval is,

\bar x - E < \mu < \bar x + E

1.4 - 0.57< \mu < 1.4+ 0.57

0.83 < \mu < 1.97

(0.83 , 1.97)